A stone falls from rest with a constant acceleration of 9.81 m/s2. How fast is it moving after 5 seconds?


1
Expert's answer
2020-09-17T09:28:20-0400

By definition, a constatn acceleration relates to the speed in the following manner:


"g = \dfrac{v}{t}"

extracting "v", obtain:

"v = gt"

Substituting numerical values: "g = 9.81m\/s^2" and "t = 5s", get:


"v = 9.81\cdot 5 = 49.05 m/s"

Answer. 49.05 m/s.